Engineering Night - UT 28 Sep 2020

OSA; J. Williams

ISp: C. Bolyard

ISA: B. Rothberg (secret bunker not on the mountai), O. Kuhn (Tucson)

Mtn Manager: J. Prothro


Night Plan:

Breakdown of Night:

  • Evening Twilight ends at 19:37 - LBC skyflats
  • 19:37-22hr LBC Engineering Tasks
    • focus/collimate
    • copoint (work with TMS system)
    • Superfocus sequence at different PAs
    • extra/intra-focal pupils
  • 22:00-22:20hr - LBC photometric standards
  • 22:20-23:30hr - 2020 SO LBC observations
  • 23:30-03:30 hr - PEPSI observations
  • 03:40-04:40 hr - GRB spectroscopic followup with MODS
  • 04:40-twlight - continue with imaging of GRB host

Possible Engineering Tasks:

  • Test SX Gregorian M1 LUT for Z9 (consumes zero time on-sky, JMH will install in afternoon)
  • LBC skyflats in twilight
  • LBC Superfocus sequence - 2-3 hrs (high priority)
    • Collect TMS data in parallel on DX
  • OSCO - can be executed by the OSA (medium priority)
  • pseudo-speckle mode (requires ESM and DL AO) - would need AO support for the night (medium priority)
  • M3 pupil wobble re-test (low priority)
  • Field aberration data collection (low priority)

Night Log

UT 00:34 - waking LBCs (Olga), and MODS (Barry)

UT 00:48 - sievesnaps complete

UT 01:30 - LBC Sky flats at Blank 19+59 field.
  • 01:40 Starting S and z flats. Counts are a bit low so I am repeating with a scale factor 5.
  • 01:48 Starting g and i flats with scale factor 3.
  • 02:00 Starting g and i flats are scaled by 7. We didn't get the LBCB flats for the 2020 SO program to be done later.

UT 02:08 - Heejoo is turning on the TMS laser

UT 02:08 = Slewed to the first field. The TMS laser is taking a while to stabilize.
  • Noting: mirrors are in pretty good equilibrium. SXg = 11.67, SXa=11.58 and DXg = 11.73 and DXa = 11.44

UT 02:25 - The TMS sw needs to be restarted so we'll collimate without it for now.

UT 02:26 - On el 70 field (67 deg), running dohybrid.
  • dohybrid on robs runs into an error when trying to write the FINDPUP_results file in /home/lbceng/WRSSave/
  • on robs, we also need to be sure we have a working directory like /scratch. I am unsure of where we can work on data on robs.
  • dohybrid from obs3. Converged but it appears that the final blue image (023951) has some residual astigmatism.
  • UT 02:43 Extra-/Intra-focal pair at +/- 0.5 (BEFORE COPOINT)
  • UT 02:47 Copointing: b 024804 and r 024800. Offsets were less than 8 arcsec
  • UT 02:54 Extra-/Intra-focal pair again (AFTER COPOINT)

UT 03:00 - On el 70 field. I am repeating the focus/collimation since it did not appear to have converged well for Blue.
  • UT 03:01 dofpia (looks like it did take out some z5 on the blue side).
  • UT 03:05 Extra-/Intra-focal (exin) pair
  • UT 03:08 superfoc PA=0
  • UT 03:17 superfoc PA=180
  • UT 03:24 superfoc PA=0
  • UT 03:32 exin focal pair (elev currently 66 deg)

UT 03:35 - Going to the el55 field at 20h30m+65
  • UT 03:36 dofpia (from robs)
    • it is not loading the ob (RB_lowextra.ob)
  • UT 03:46 dofpia finished (had to run it on obs3)
  • UT 03:46 copointing b 034645 and r 034640. Corrections were <~2.6" on the mirror and on the mount, also delta CE < 3"
  • UT 03:52 Switching over to the winter chiller
  • UT 03:53 exin focal pair
  • UT 03:56 superfoc PA=0
  • UT 04:07 superfoc PA=180
  • UT 04:13 superoc PA=0
  • UT 04:22 exin focal pair

UT 04:22 The TMS laser is running. It was necessary to powering off/on the TMS machine in the AUX control room.

UT 04:25 - Going to el75a field at 23h +35d dec.
  • UT 04:27 running dofpia
  • UT 04:33 copointing b 43418 and 43414. Mirror motions are < 1.5" but mount offset is larger, CE ~ 11".
    • This was the largest copointing correction of the night, larger than the starting one.
  • UT 04:38 exin focal pair
  • UT 04:42 superfoc PA=0
  • UT 04:50 superfoc PA=180
    • Heejoo has noted that as M1 is commanded to step in Z, he sees a change in the diameter along the elevation axis (z6). The diameter in the other direction does not change.
  • UT 05:05 exin focal pair (at 80 deg elevation - this is about as high as I am comfortable going, since we have seen trailing at el>80).
  • UT 05:11 I am taking a pair of 100-sec in-focus images at this field which just rose above 80-deg elevation, to see if there is evidence of rotational trailing and also to get a pair of in-focus images. I took 2 pairs, because the first one used only chip2 (I just scaled the exposure time of the copointing OB - I edited it to use the 4 chips).

UT 05:27 Slewed to copointing field for 2020SO.
  • UT 05:27 running dofpia here. Elevation is 38 deg. FPIA estimates seeing 1.50"
  • UT 05:34 copointing b 053400 r 053356. delta CD is ~7", mirror corrections are <3".
  • UT 05:37 running the OB.
    • Seeing on the guiders is 1.4" and 1.2".
    • Background counts are pretty high (17k on the blue and 12k on the red). These levels are ok, but higher than I had expected. The target is 60 deg from the moon.
    • We see the object by blinking a pair of g-SLOAN exposures. At g-SLOAN, the peak is about 500 counts above the background.
    • At UT 06 (elev 44), the focus on the guiders is better - 1"/0.8" on blue/red.
    • In i_SLOAN (z_SLOAN), the object has ~3000 (1500) ADU above the background.

UT 06:13 I am running the OB again. The focus/collimation looked pretty good, judging from the guide star FWHM values, and we should hand over to PEPSI soon.
  • UT 06:15 FWHM is 1.2", 1.1" on guiders.
  • UT 06:32 Stopped the sequence to reconfigure for PEPSI.

UT 06:47 - still stuck on authorizing, waiting for swing arms to finish move before trying anything

UT 07:14 - still working on SX authorization issue. J. Williams talking Y.Zhang.

UT 07:19 - still working on authorization issue.

UT 07:22 - AZcam server for SX for PEPSI had hung, restarting.

-now reauthorizing again

UT 07:24 -- success

UT 07:35 - acquiring poitnting stars - some issue with focus being off, now fixed

UT 07:49 - starting science for first PEPSI target

UT 08:10 - slewing to 2nd target

UT 09:03 -slewing to 4th target

UT 10:09 - moving to final target

UT 10:33 - finished with PEPSI, reconfigruing for Bino-MODS

UT 10:49 - pointing check

-forgot to open hatch, all ok now

UT 11:03 sending binary acquisition script

UT 11:12 - 60 sec cant see, trying 120sec if can't see it then , will switch to imaging

UT 11:16 - can't see it in 120sec switching to imaging g and r

-MODS1 there's a huge glint in MODS1 that seems to affect the background, i see a smudge in MODS-2, although it doesn't quite look in the right place.

UT 11:18 - running 180sec exposures in g and r-sloan with 4 pt dither (+/-15")

UT 11:40 - can stil see it faintly in all channels seeing ranging from 1-1.2" but a bit more variable than earlier

UT 11:55 - MODS2R 0011 the stars all jumped - DO NOT USE

-WFS/Guider showing 1" at airmass 1.9!

UT 12:10 - hitting 30 elevation limit, letting MODS-2 readout now then stopping

M2B 002-0012, M2R 005-0015

M1B 0002-0012, M1R 005-0015

UT 12:14 - attempting to do skyflats for MODS using blank field 04 +25

-MODS1 got stuck ---trying again

UT 12:25 - trying 10sec exposures ... probably still too dark

idling for skyflats

UT 12:30 trying agan

trying 20sec

M1B 0017 M1R 0020 ; M2B 0017, M2R 0019 (offsets -30,-30)

5-6k for MODS2, 3-4k for MODS1

M1B 0018, M1R 0021, M2B 0018, M2R 0020 (offsets -30,-30) counts 5-6k, 8-9l

M1B 0019, M1R 0022, M2B 0019, M2R 0021 (offsets -20,-20) counts 8-9k, 11-12k

M1B 0020, M1R 0023, M2B 0020, M2R 0022 (offsets -20,-10), 12-17j, 21-24k

M1B 0021, M1R 0024, M2B 0021, M2R 0023 (offsets -10, -10, exptime 10seconds) ~20k counts

M1B 0022, M1R 0025, M2B 0022, M2R 0024 (offsets 0,0, exptime 5seconds) 8-12k, 15-16k

M1B 0023, M1R 0026, M2B 0023, M2R 0025 (offsets 5,5 exptime 5 seconds) 13-20k, 23k

M1B 0024, M1R 0027, M2B 0024, M2R 0026 (offsets 10,10 exptime 3 seconds) 25k

UT 12:48 - closing up, got decent skyflats

Calibrations g and r internal flats

M1B 0025-0029

M1R 0028-0032

M2B 0025-0029

M2R 0027-0031

Bias 3Kx3K

M1B 0030-0034

M1R 0033-0037

M2B 0030-0034

M2R 0032-0056

UT 13:25 - putting both MODS to sleep

-- BarryRothberg - 22 Sep 2020
Topic revision: r11 - 28 Sep 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback